Hey guys, I am having a lot of trouble with my homework. Could you please help? I thumbs up anyone who gives a thorough explanation and shows the work! I really need to understand these concepts!1) At 25 °C, the pH of a 0.15 M aqueous solution of CsZ (the cesium salt of HZ) is 10.70. What is the Ka for HZ? 2) Calculate

Thank you!
Answer #1

1)

given, pH = 10.70 \Rightarrow pOH = 14 - pH = 3.30 \Rightarrow [OH-] = 5.01 \times 10-4 M.

the reaction is CsZ + H2O  \rightleftharpoons  Cs+ + OH- + HZ

initial concentrations: 0.15 0 0 0

final concentrations: 0.15-x x x x

x =  5.01 \times 10-4 M << 0.15M.

Therefore, 0.15-x = 0.15(approximately).

Now, [Z-](which is nothing but the [CsZ] because CsZ exists as Cs+ and Z- in water) = 0.15 <.

[HZ] = x = 5.01 \times 10-4 M.

[H+] = 2 \times 10-11 M

The dissociation reaction of acid HZ is: HZ \rightleftharpoons    H+ + Z-

Ka = [H+][Z-]\div[HZ] = ( 2 \times 10-11 M ) \times (0.15 M ) \div ( 5.01 \times 10-4 M ) = 6.0  \times 10-9 M

So, we just have to find the concentration and us appropriate equation to find the dissociation constants.

We can only answer one question.